# Peer-induced Fairness: A Causal Approach for Algorithmic Fairness Auditing

### v3

@@ Abstract
As artificial intelligence and automation increasingly permeate decision-making systems, ensuring algorithmic fairness has become critical. This paper addresses a fundamental question often overlooked: how to audit algorithmic fairness scientifically. It is crucial to discern whether adverse decisions stem from algorithmic discrimination or merely from the subjects’ insufficient capabilities. To tackle this, we develop an algorithmic fairness auditing framework, “peer-induced fairness”, that leverages counterfactual fairness and advanced causal inference techniques, such as the Single World Intervention Graph. Our approach transcends the typical trade-off between quantitative fairness measures and accuracy by aiming to assess algorithmic fairness at the individual level through peer comparisons and hypothesis testing, particularly in contexts like credit approval. This framework effectively addresses data scarcity and imbalance—frequent data quality challenges in traditional models, and is a uniquely model-agnostic and flexible self-audit tool for stakeholders and an external audit tool for regulators, in a plug-and-play fashion. Additionally, it offers explainable feedback for those who receive unfavourable decisions due to insufficient capabilities. We validate our framework in a practical context, highlighting the degree of algorithmic bias that arises, with 41.51% and 56.40% of subjects being either discriminated against or privileged respectively. It could also serve as a transparent, and adaptable tool suitable for diverse applications.
Key Words: Ethics in OR, Algorithmic Fairness, Causality, Counterfactual Fairness
@@ 1 Introduction
Algorithmic data-driven methods are extensively employed across a broad spectrum of fields, including healthcare, advertising, employment, supply chain, credit scoring, criminal justice (Kozodoi et al.,, 2022; Guan et al.,, 2020; Chen and Hooker,, 2022; Berk et al.,, 2017; Dwork et al.,, 2012; Lodi et al.,, 2023, 2024). These approaches are being adopted to replace human decision-making with the aim of reducing biases and ostensibly moving society towards greater equality. This is because algorithms and robots, as non-human entities, do not inherently possess the biases that can influence human judgement, thereby reducing the potential for discrimination. Admittedly, algorithms do not discriminate against groups of individuals; however, algorithmic data-driven models often rely on historical datasets that may contain significant biases. There are a number of potential areas of use of these datasets where algorithmic bias might arise. Therefore, to truly achieve the objective of reducing discrimination, it is crucial to develop algorithmic models that are specifically designed to handle and correct for biases within these datasets. Such models must effectively make decisions that are free from discrimination based on protected characteristics, such as gender or marital status (Lessmann et al.,, 2015; Kozodoi et al.,, 2022).
Such algorithmic fairness is not only applicable to people but exists in any context where an organisation may be treated unfairly in a decision process. As regulatory policies and documents, such as the General Data Protection Regulation (GDPR) (Voigt and Von Dem Bussche,, 2017) and the Markets in Financial Instruments Directive II (MiFID II) (Yeoh,, 2019), continue to evolve, algorithmic fairness auditing and monitoring becomes increasingly significant. The Europe Artificial Intelligence Act (Madiega,, 2021) (EU AI Act), the first comprehensive AI law in the world, requires that high-risk applications, such as credit scoring, to identify discrimination themselves. It also mandates that AI systems in these applications should be assessed before being put on the market and continuously throughout their life cycle. British Standards Institution (BSI) subsequently requires that AI providers ensure full compliance with the EU AI Act (British Standards Institution,, 2023). To support clients who will be regulated by this legislation, BSI emphasises the significance of algorithm auditing services, which are designed to help AI providers meet the necessary standards and regulatory requirements set by the EU AI Act (British Standards Institution,, 2023). Therefore, offering a reasonable tool for self-audit for stakeholders, as well as for external audits by regulators, is crucial.
However, a key challenge is obtaining precise and stable auditing results, especially given poor data quality, such as data scarcity and imbalance, which are common issues in many types of research. Additionally, it is crucial to distinguish whether the rejection of an individual or organisation is due to discrimination or inherent incapability.
In response, our paper employs a straightforward yet effective causal framework to audit the existence of algorithmic bias—that is, to determine whether certain individuals or organisations are being unfairly treated. The core concept involves comparing treatment in model outcomes among similarly situated individuals or organisations within the dataset. By analysing these comparisons, we aim to ascertain whether any individual or organisation is being unfairly treated by the algorithm.
Although the concept of comparing similar individuals is neither new nor complex, our framework, termed “peer-induced fairness” makes significant contributions to the field. First, this is the first framework to formalise a practical concept of “peer-induced fairness” specifically designed to audit algorithmic biases. Unlike traditional static measures of fairness, “peer-induced fairness” is an advanced framework that leverages counterfactual fairness (Wu et al.,, 2019) and causal inference techniques, such as Single World Intervention Graphs (SWIGs) (Richardson and Robins,, 2013) and peer observation theory (Li and Jain,, 2016; Ho and Su,, 2009). Stakeholders and regulators could use it as a bias audit tool for self-assessment and external assessment. Second, the core counterfactual comparison approach makes “peer-induced fairness” robust against data scarcity issues—a common challenge where protected groups are often underrepresented in datasets. This methodology circumvents the need for traditional statistical estimates within the protected group by utilising robust counterfactual statistics derived from well-represented peer groups. This approach allows us to use data from a single group, thereby making our method robust to population imbalance. Third, “peer-induced fairness” offers a transparent framework that could distinguish the subjects who are unfairly treated and merely not capable. It enables comparisons between an individual’s and their peers’ data, providing watch-out insights regarding the key features of why they are fairly treated but still rejected. Fourth, we validate our framework on access to finance for small and medium-sized enterprises (SMEs). Our application highlights in detail many advantages of using the framework which details unfairness and other aspects such as explainability. Given it relates to organisations, it expands the literature from predominantly individual-focused studies to include corporate entities, using firm size as a protected characteristic, given that smaller businesses are more likely to be denied loans and unfairly treated in accessing finance. Fifth, our criteria and framework are adaptable, offering potential for broader application across various fields and datasets. It is not only applicable to credit scoring for individuals but also extends to firms and can be adapted to any domain requiring fairness analysis. Its versatility stems from its ability to handle multifaceted scenarios independently of the specific fairness measures or the underlying problem because it assesses peers in a counterfactual manner.
The rest of our paper is structured as follows. Section 2 reviews the relevant literature. Section 3 starts from the counterfactual world and introduces the casual framework. Section 4 proposes the peer observation theory, corresponding peer identification process and “peer-induced fairness” framework. Section 5 and Section 6 demonstrate our experiment procedure and empirical results. Section 7 concludes.
2 Literature review
The theory and practice surrounding fairness have garnered increasing attention from both scholars and regulators (Federal Trade Commission,, 2023; Rohner,, 1979; Voigt and Von Dem Bussche,, 2017; Kehrenberg et al.,, 2020). The concept of algorithmic fairness in automated decision-making systems is notably complex and lacks a universally accepted definition. Several frameworks have been proposed to address this challenge (Dwork et al.,, 2012; Hardt et al.,, 2016).
Despite advancements in developing measures to uphold fairness criteria, there remains a considerable gap in the practical application of these frameworks. Traditionally, academic efforts have focused on transforming the concept of fairness into quantifiable definitions that address discrimination within specific datasets. However, those responsible for implementing these algorithms—such as practitioners, policymakers, and judicial figures—face significant challenges in choosing the most appropriate fairness definition to suit their unique circumstances (Kusner et al.,, 2017; Huang et al.,, 2020; Dixon et al.,, 2018; Foulds et al.,, 2020; Hickey et al.,, 2020). For example, the criteria for fairness required to address gender disparities may vary markedly from those needed for racial issues, and similarly, from the broader, non-demographic contexts, such as ensuring equitable treatment between large corporations and SMEs in credit approval processes (Lu and Calabrese,, 2023). It is impractical to adopt a single quantitative fairness definition as a universal solution for all sectors. Furthermore, it is essential to recognise that, despite their positive intentions, some fairness models can inadvertently increase discrimination (Kozodoi et al.,, 2022). This highlights the urgent need for a detailed and context-specific evaluation of fairness definitions, to ensure that the deployment of algorithmic decision-making systems genuinely contributes to reducing bias (Kusner et al.,, 2017).
In response to persistent issues in algorithmic decision-making, a causally-oriented approach to fairness has been advocated (Kusner et al.,, 2017), focusing on the relationships between protected features and data. Subsequent studies (Pfohl et al.,, 2019; Kim et al.,, 2021; Kusner et al.,, 2017; Chiappa,, 2019) have shown the efficacy of causal inference techniques in developing fair algorithms. However, counterfactual fairness encounters significant limitations, such as unidentifiability from observational data under certain conditions, which complicates the measurement of counterfactual outcomes (Wu et al.,, 2019). Additionally, the challenge of data scarcity often hinders decision-making processes intended to implement fairness constraints. Historical biases typically result in datasets where protected groups are underrepresented (Iosifidis and Ntoutsi,, 2018), thereby skewing the accuracy of fairness criteria and utility metrics. This imbalance, particularly the under-representation of minority groups in training data (i.e., representational disparity), leads to their diminished influence on model objectives (Hashimoto et al.,, 2018). As a result, biased measures of discrimination may emerge (Sha et al.,, 2023; Dablain et al.,, 2022). For example, in the finance sector, the availability of credit approval data for minority groups is substantially lower than for majority groups, complicating the fair assessment of creditworthiness—a critical aspect of many established fairness frameworks. Furthermore, the implementation of complex causal frameworks often depends on intricate causal graph assumptions and elaborate causal inferences, such as those detailed in (Chiappa,, 2019). These calculations are not only complex but also typically require Monte Carlo approximations. While these frameworks are adequate for group-level fairness analyses, they are less suited to addressing the needs of specific individuals or firms. Those seeking to enhance their chances of approval for future financing applications require more tangible explanations and actionable feedback than that which Monte Carlo approximations can provide. Regulatory authorities have consistently emphasised the necessity for transparent and explainable models to provide clear decision-making grounds (Chen et al.,, 2024; Voigt and Von Dem Bussche,, 2017). However, current explanation-related fairness criteria usually incorporate explainability into the fairness framework (Zhao et al.,, 2023; Hickey et al.,, 2020). There is still a gap in providing explanations of the fairness framework, which is crucial as it aids people in understanding the specific reasons behind the rejections. Therefore, this paper contributes to filling the gaps by designing a novel fairness framework using a causal lens, to stably audit algorithmic bias with group imbalance and data scarcity, and provide explanations to promote the transparency of our framework.
3 Counterfactual fairness and SWIGs
The attractiveness of counterfactual reasoning stems from its capacity to rigorously analyse causal relationships, unearth potential biases, and furnish methodologies for elucidating decisions made by models. Counterfactual reasoning critically examines and establishes causal connections by contemplating hypothetical scenarios under altered conditions (e.g., “If the individual were not a woman, would her application be approved for a loan?”). Counterfactual fairness is a concept that has been explored and represented in diverse forms within the academic literature (Pfohl et al.,, 2019; Kim et al.,, 2021; Kusner et al.,, 2017; Wu et al.,, 2019). In this paper, we adopt the general framework as described by Wu et al., (2019).
@@ 3 Counterfactual fairness and SWIGs
For illustrative purposes, imagine a scenario where individuals/organisations are evaluated for accessing finance using a predictive model, which determines the decision outcome, represented as $\hat{Y}$. Let us focus on a firm from the smallest SME (i.e., micro group), denoted by $s_{-}$ with a specific profile $\bm{x}$. The likelihood that this firm receives a favourable outcome is expressed as $\mathbb{P}(\hat{Y}|s_{-},\bm{x})$, which is equivalent to $\mathbb{P}(\hat{Y}_{S\leftarrow s_{-}}=1|S=s_{-},\bm{X}=\bm{x})$ by maintaining the firm’s protected feature (i.e., firms’ size) unaltered. Suppose, hypothetically, that this firm’s protected feature is changed from $s_{-}$ to $s_{+}$. The probability of a favourable outcome after such a counterfactual modification is denoted by $\mathbb{P}(\hat{Y}_{s_{+}}|s_{-},\bm{x})$. Counterfactual fairness is achieved when the probabilities $\mathbb{P}(\hat{Y}_{s_{-}}|s_{-},\bm{x})$ and $\mathbb{P}(\hat{Y}_{s_{+}}|s_{-},\bm{x})$ are equal, suggesting that the treatment of the firm would remain consistent irrespective of their group membership. This condition underscores the essence of counterfactual fairness, where the decision-making process is indifferent to changes in the protected features of the firms.
@@
Figure 1: SWIGs for Graphical Causal Models (GCM). (a): The SWIG $\mathcal{G}(s_{-},\bm{x})$ represents the actual scenario for an individual with features $(s_{-},\bm{x})$. (b): The SWIG $\mathcal{\tilde{G}}(s_{+},\bm{x})$ illustrates the counterfactual scenario, assuming the individual’s protected feature changes from $s_{-}$ to $s_{+}$, while their other features $\bm{x}$ remain the same. (c): The SWIG $\mathcal{G}(s_{+},\bm{x}^{\prime})$ represents the actual scenario for an individual with features $(s_{+},\bm{x}^{\prime})$. The actual SWIG $\mathcal{G}(s_{-},\bm{x})$ corresponds to the conditional distribution $\hat{Y}_{s_{-}}|s_{-},\bm{x}$. Conversely, in the counterfactual SWIG $\mathcal{\tilde{G}}(s_{+},\bm{x})$ refers to $\hat{Y}_{s_{+}}|s_{-},\bm{x}$, denoting the outcome distribution had the individual been featured with $s_{+}$, given that the actual features are $(s_{-},\bm{x})$. Thus the directed link from $s_{+}$ to $\bm{X}(s_{-})$ is not the fact (shown in green colour). Note: $\mathcal{\tilde{G}}(s_{+},\bm{x})\neq\mathcal{G}(s_{+},\bm{x}^{\prime})$ because $\mathcal{\tilde{G}}(s_{+},\bm{x})$ is counterfactual scenario with actual features $(s_{-},\bm{x})$ and $\mathcal{G}(s_{+},\bm{x}^{\prime})$ is the fact with features $(s_{+},\bm{x}^{\prime})$.
A more nuanced comprehension of counterfactual fairness may be facilitated through the lens of SWIGs [^1] In SWIGs, black nodes represent random variables, while red nodes indicate fixed values, representing experimental interventions. Arrows depict causal relationships between variables. (Richardson and Robins,, 2013). Consider an individual belonging to a disadvantaged group $s_{-}$, characterised by features $\bm{x}$. The label $s_{-}$ could exert a direct influence on the outcome $Y$, or it may indirectly impact $Y$ through its effect on other observable features $\bm{X}$. If we postulate a counterfactual scenario in which the individual’s group designation changes from $s_{-}$ to $s_{+}$, the corresponding Graphical Causal Models (GCMs) for both actual and hypothetical situations can be depicted using SWIGs, as illustrated in Fig. 1. Counterfactual fairness is attained if the predictor, consistent with the actual GCM and the counterfactual GCM, yields identical probabilities for the outcome given the specific features $(s_{-},\bm{x})$.
Next, let us review some pivotal conclusions derived from the SWIGs as depicted in Panel (a) of Fig. 1 and propose some notations. A key aspect we will discuss is the factorisation properties of the joint distribution of all variables within a SWIG, applicable to any protected feature $s_{-},s_{+}$ and other features $\bm{x}$, which can be mathematically represented as follows:
@@
highlighting the left-hand side is the potential outcome while the right-hand side is the observational conditional probability. In the context of the counterfactual scenario with actual features $(s_{-},\bm{x})$ shown in Panel (b) of Fig. 1, a similar joint distribution is applicable:
@@
4 Peer-induced fairness with causal framework
While the concept of counterfactual fairness is theoretically straightforward and can be easily described, its application in practice is hampered by the challenges in identifying counterfactual outcomes from observational data in certain scenarios, as highlighted by Wu et al., (2019). Specifically, the probability $\mathbb{P}(\hat{Y}_{s_{+}}|s_{-},\bm{x})$ as a potential outcome remains elusive for direct calculation due to its unidentifiability. To navigate this impediment and facilitate a feasible implementation of counterfactual fairness, we propose a practical approximation method that utilises peer comparison as an effective strategy.

### v4

@@ Abstract
With the European Union’s Artificial Intelligence Act taking effect on 1 August 2024, high-risk AI applications must adhere to stringent transparency and fairness standards. This paper addresses a crucial question: how can we scientifically audit algorithmic fairness? Current methods typically remain at the basic detection stage of auditing, without accounting for more complex scenarios. We propose a novel framework, “peer-induced fairness”, which combines the strengths of counterfactual fairness and peer comparison strategy, creating a reliable and robust tool for auditing algorithmic fairness. Our framework is universal, adaptable to various domains, and capable of handling different levels of data quality, including skewed distributions. Moreover, it can distinguish whether adverse decisions result from algorithmic discrimination or inherent limitations of the subjects, thereby enhancing transparency. This framework can serve as both a self-assessment tool for AI developers and an external assessment tool for auditors to ensure compliance with the EU AI Act. We demonstrate its utility in small and medium-sized enterprises’ access to finance, uncovering significant unfairness—41.51% of micro-firms face discrimination compared to non-micro firms. These findings highlight the framework’s potential for broader applications in ensuring equitable AI-driven decision-making.
Key Words: Algorithmic fairness, AI auditing, Causality, Counterfactual fairness, Small and medium-sized enterprises, Credit scoring
@@ 1 Introduction
The increasing adoption of data-driven methodologies across various sectors has heightened global concerns about algorithmic bias. These concerns are underscored by the recent enactment of the European Union’s Artificial Intelligence Act (EU AI Act), effective from 1 August 2024 (Madiega,, 2024). The Act represents the world’s first comprehensive legal framework for Artificial Intelligence (AI). It imposes stringent requirements on high-risk AI applications, such as credit scoring systems, mandating the rigorous identification and mitigation of discrimination risks. Additionally, the Act requires that these AI systems undergo thorough assessments both prior to deployment and continuously throughout their operational life cycle, thereby ensuring sustained compliance with transparency and fairness standards.
The implementation of the EU AI Act necessitates the urgent development of a universal and transparent tool capable of ensuring compliance with rigorous standards. This task is critical not only for regulators but for everyone involved in the AI community. A key challenge lies in scientifically assessing the fairness of decisions made by AI models—a question that remains at the forefront of ethical AI development.
Over the past decade, the importance of algorithmic fairness has increasingly gained recognition as a vital component of responsible technology use and ethical AI development. Despite this growing awareness, a substantial body of literature highlights the trade-offs between accuracy and fairness (Kim et al.,, 2023; Huang et al.,, 2020; Guldogan et al.,, 2022; Dixon et al.,, 2018; Foulds et al.,, 2020; Chen et al.,, 2022; Hickey et al.,, 2020; Dwork et al.,, 2012; Hardt et al.,, 2016). However, there remains a significant gap in the development of robust auditing frameworks specifically designed to assess and monitor algorithmic bias. Some studies have proposed approaches for auditing algorithmic fairness with data-driven models (Cherian and Candès,, 2024; Brundage et al.,, 2020; Xue et al.,, 2020; Tramer et al.,, 2017; Si et al.,, 2021; Yan and Zhang,, 2022). However, these approaches primarily focus on fairness detection based on internally accessible data and models. When regulators obtain data externally, the underlying algorithmic models and decision-making processes are often unknown, necessitating a reassessment of whether the decisions produced by AI developers’ models are fair. Besides, although identifying the presence of bias is necessary, it represents only the initial step in the comprehensive auditing of AI systems, and it is insufficient for ensuring robust and reliable auditing frameworks. The two key factors concerning the complexities inherent in the real world—universality and transparency—are often overlooked.
With respect to universality, a mature auditing framework should be adaptable to datasets with different characteristics. For example, a critical challenge in the development of auditing tools is the issue of data quality, particularly data scarcity and imbalance, which are pervasive in real-world datasets (Lessmann et al.,, 2015; Chen et al.,, 2024; Sha et al.,, 2023; Dablain et al.,, 2022). Historical biases frequently result in the under-representation of protected groups within datasets (Iosifidis and Ntoutsi,, 2018). This imbalance, particularly the under-representation of minority groups in training data (i.e., representational disparity), leads to their diminished influence on model objectives and reduces their influence on model objectives (Hashimoto et al.,, 2018). Consequently, biased discrimination measures and unreliable audit outcomes may emerge (Sha et al.,, 2023; Dablain et al.,, 2022; Sha et al.,, 2022; Yan et al.,, 2020). A particularly concerning issue arises when an individual or organisation is flagged as discriminated against in one dataset but deemed privileged in another due to varying degrees of imbalance. Such inconsistencies pose significant challenges to conducting universal audits. Current frameworks frequently assume that data is balanced across different groups, an assumption rarely met in practice. Some studies propose oversampling techniques, such as SMOTE, to address data imbalance (Sha et al.,, 2023, 2022; Yan et al.,, 2020). However, these methods can inadvertently “smooth out” critical features within the data, thereby altering the intricate relationships between variables (Chen et al.,, 2024), which may ultimately lead to distorted auditing outcomes. Such flawed outcomes can be disastrous for regulators, companies, and third-party auditors, as they could introduce even greater risks. Audits based on unrealistic assumptions may fail to detect genuine instances of algorithmic discrimination, allowing unfair practices to persist. Furthermore, inaccurate audits could lead to misguided adjustments in algorithms, potentially worsening performance or introducing new biases, thus compounding the original issues. If stakeholders, including the public, perceive the auditing process as flawed or unreliable, it can erode confidence in both the regulators and the entities being audited. In terms of transparency, both regulatory authorities and researchers have consistently underscored the importance of transparent and explainable models that provide clear justifications for decision-making (Chen et al.,, 2024; Voigt and Von Dem Bussche,, 2017). Current fairness-related criteria often incorporate explainability into the fairness assessment (Zhao et al.,, 2023; Hickey et al.,, 2020; Chen et al.,, 2022). However, a gap remains in delivering clear and understandable explanations. This gap is critical, as understanding the reasons behind rejections is essential for ensuring that decisions are perceived as fair. An opaque auditing framework can similarly lead both the audited entities and the public to perceive the audit process as incredible, thereby undermining its credibility and public confidence. Therefore, an effective auditing framework must elucidate the reasons for adverse decisions, particularly to distinguish whether such decisions are due to discrimination or inherent limitations.
We propose a reliable and robust audit framework that solves the above issues, embodying both universality and transparency. It is a causally-oriented approach to fairness (Kusner et al.,, 2017). Compared to traditional static fairness criteria, a causally-oriented approach offers a more effective means of addressing real-world challenges. This approach is particularly valuable for practitioners, policymakers, and judicial authorities tasked with implementing algorithms designed to mitigate discrimination. Selecting an appropriate fairness definition that aligns with the specific nuances of each scenario can be a significant challenge. For instance, the parameters for fairness in addressing gender disparities may differ substantially from those relevant to racial issues, or from considerations in broader, non-demographic contexts, such as ensuring equitable treatment between large corporations and SMEs in credit approval processes. Thus, applying a single quantitative fairness definition as a universal remedy across all sectors is impractical. This reasoning underpins our reliance on the causal framework, previous studies (Gastwirth,, 1997; Pfohl et al.,, 2019; Kim et al.,, 2021; Kusner et al.,, 2017; Chiappa,, 2019; Imai et al.,, 2023, 2013) have demonstrated the efficacy of causal inference techniques, with counterfactual causal inference standing out as particularly prominent. Counterfactual reasoning critically examines and establishes causal connections by contemplating hypothetical scenarios under altered conditions (e.g., “If the applicant were not a female, would the application be approved for a loan?”). While counterfactual approaches to fairness have been previously suggested (Kusner et al.,, 2017), a key limitation of counterfactual fairness is the unidentifiability from observational data (Wu et al.,, 2019). To address this, our framework proves that individuals or organisations with similar joint distributions could be identified as counterfactual instances. We identify these counterfactual instances as peers. We are motivated by the peer comparison perception (Li and Jain,, 2016), which involves the differential treatment experienced by individuals compared to their peers. When an individual’s treatment is consistent with that of their peer group, perceptions of bias tend to diminish.
Building on this insight, we introduce a novel concept termed “peer-induced fairness”, which leverages the strengths of counterfactual fairness while addressing its limitations, thereby creating a more reliable and robust tool for auditing algorithmic fairness. This framework also serves as a valuable self-assessment tool, which is increasingly crucial for the AI community in the development of products that must comply with the EU AI Act. Beyond its general contributions to the field of algorithmic fairness auditing and self-assessment, our paper has the following particular contributions: First, to the best of our knowledge, our framework is the first to formalise a practical concept of “peer-induced fairness” specifically designed to audit algorithmic biases. This comprehensive framework goes beyond the initial stage of basic detection, enabling users to evaluate externally obtained data without accessing the decision-making process or the underlying algorithmic model. Second, our framework is universal and versatile in handling different levels of data quality, including datasets with highly skewed distributions of protected attributes issues often overlooked or inadequately addressed in previous studies. Third, when it is considered as a self-assessment tool, the “peer-induced fairness” framework not only provides conclusions from self-assessment but also offers insightful explanations through peer comparisons, enhancing transparency and explainability. Fourth, we demonstrate the practical utility of the “peer-induced fairness” framework in uncovering algorithmic fairness issues related to small and medium-sized enterprises (SMEs) access to finance, particularly in scenarios where AI systems are used for decision-making. To the best of our knowledge, this is the first study of algorithmic bias on SMEs’ access to finance [^1] Lu and Calabrese, (2023) proposed a method for assessing the discrimination in ground truth $Y$ in SMEs’ access to finance, rather than in the algorithmic predictions ($\hat{Y}$) made by AI systems. It also highlights the effectiveness of our framework as a self-assessment tool in real-world applications. Besides, our framework is universal, applicable across different domains and capable of addressing various types of protected attributes. For example, it expands the protected attribute from the customer level (e.g., gender and race) to the organisation level (i.e., firm size).
The remainder of this paper is organised as follows. Section 2 begins with an overview of the foundational concepts of counterfactual fairness and causal framework, including the representation of Single World Intervention Graphs (SWIGs). In Section 3, we introduce our “peer-induced fairness” framework in a step-by-step manner. Sections 4 and 5 detail our experimental procedures and present the empirical results, using the example of SMEs’ access to finance in the UK. Finally, concluding remarks and further discussions are provided in Section 6.
2 Counterfactual fairness and its representation
Before presenting our framework, it is essential to introduce some key concepts related to counterfactual fairness and its representation. Various forms of counterfactual fairness have been proposed in the academic literature (Pfohl et al.,, 2019; Kim et al.,, 2021; Kusner et al.,, 2017; Wu et al.,, 2019). In this paper, we adopt the general framework outlined by Wu et al., (2019).
@@ 2 Counterfactual fairness and its representation
For illustrative purposes, imagine a scenario - loan applications using a predictive model, which determines the decision outcome, represented as $\hat{Y}$. Let us focus on an application by a female, denoted by $s_{-}$ with a specific profile $\bm{x}$. The likelihood that this applicant received a favourable outcome is expressed as $\mathbb{P}(\hat{Y}|s_{-},\bm{x})$, which is equivalent to $\mathbb{P}(\hat{Y}_{S\leftarrow s_{-}}=1|S=s_{-},\bm{X}=\bm{x})$ by maintaining the protected feature (i.e., gender) unaltered. Suppose, hypothetically, that this applicant’s protected feature is changed from $s_{-}$ to $s_{+}$. The probability of a favourable outcome after such a counterfactual modification is denoted by $\mathbb{P}(\hat{Y}_{s_{+}}|s_{-},\bm{x})$. Counterfactual fairness is achieved when the probabilities $\mathbb{P}(\hat{Y}_{s_{-}}|s_{-},\bm{x})$ and $\mathbb{P}(\hat{Y}_{s_{+}}|s_{-},\bm{x})$ are equal, suggesting that the treatment of the applicant would remain consistent irrespective of their membership. This condition underscores the essence of counterfactual fairness, where the decision-making process is indifferent to changes in the protected features.
A more nuanced comprehension of counterfactual fairness may be facilitated through the lens of SWIGs (Richardson and Robins,, 2013). Consider an individual belonging to a disadvantaged group $s_{-}$, characterised by features $\bm{x}$. The label $s_{-}$ could exert a direct influence on the outcome $Y$, or it may indirectly impact $Y$ through its effect on other observable features $\bm{X}$. If we postulate a counterfactual scenario in which the individual’s group designation changes from $s_{-}$ to $s_{+}$, the corresponding Graphical Causal Models (GCMs) for both actual and hypothetical situations can be depicted using SWIGs, as illustrated in Figure 1. Counterfactual fairness is attained if the predictor, consistent with the actual GCM and the counterfactual GCM, yields identical probabilities for the outcome given the specific features $(s_{-},\bm{x})$.
@@
Figure 1: SWIGs for Graphical Causal Models (GCM). The nodes with black border represent random variables, while red ones indicate fixed values of random variables, representing experimental interventions. Arrows depict causal relationships between variables. (a): The SWIG $\mathcal{G}(s_{-},\bm{x})$ represents the actual scenario for an individual with features $(s_{-},\bm{x})$. (b): The SWIG $\mathcal{\tilde{G}}(s_{+},\bm{x})$ illustrates the counterfactual scenario, assuming the individual’s protected feature changes from $s_{-}$ to $s_{+}$, while their other features $\bm{x}$ remain the same. (c): The SWIG $\mathcal{G}(s_{+},\bm{x}^{\prime})$ represents the actual scenario for an individual with features $(s_{+},\bm{x}^{\prime})$. The actual SWIG $\mathcal{G}(s_{-},\bm{x})$ corresponds to the conditional distribution $\hat{Y}_{s_{-}}|s_{-},\bm{x}$. Conversely, in the counterfactual SWIG $\mathcal{\tilde{G}}(s_{+},\bm{x})$ refers to $\hat{Y}_{s_{+}}|s_{-},\bm{x}$, denoting the outcome distribution had the individual been featured with $s_{+}$, given that the actual features are $(s_{-},\bm{x})$. Thus the directed link from $s_{+}$ to $\bm{X}(s_{-})$ is not the fact (shown in green colour). Note: $\mathcal{\tilde{G}}(s_{+},\bm{x})\neq\mathcal{G}(s_{+},\bm{x}^{\prime})$ because $\mathcal{\tilde{G}}(s_{+},\bm{x})$ is counterfactual scenario with actual features $(s_{-},\bm{x})$ and $\mathcal{G}(s_{+},\bm{x}^{\prime})$ is the fact with features $(s_{+},\bm{x}^{\prime})$.
Next, let us review some pivotal conclusions derived from the SWIGs as depicted in Figure 1 (a) and propose some notations. A key aspect we will discuss is the factorisation properties of the joint distribution of all variables within a SWIG, applicable to any protected feature $s_{-},s_{+}$ and other features $\bm{x}$, which can be mathematically represented as follows:
@@
highlighting the left-hand side is the potential outcome while the right-hand side is the observational conditional probability. In the context of the counterfactual scenario with actual features $(s_{-},\bm{x})$ shown in Figure 1 (b), a similar joint distribution is applicable:
@@
While the above concept of counterfactual fairness is theoretically straightforward and can be easily described, its application in practice is hampered by the challenges in identifying counterfactual outcomes from observational data in certain scenarios, as highlighted by Wu et al., (2019). Specifically, the probability $\mathbb{P}(\hat{Y}_{s_{+}}|s_{-},\bm{x})$ as a potential outcome remains elusive for direct calculation due to its unidentifiability. This creates a significant challenge for regulators, making it difficult to implement algorithmic bias auditing.
3 Peer-induced fairness with causal method
In this section, we propose a practical approximation method that utilises peer comparison as an effective strategy, in order to navigate the above impediment and facilitate a feasible implementation of counterfactual fairness. We then introduce our “peer-induced fairness” framework and algorithmic bias-detecting methods after providing the peer identification concept.